OpenStudy (anonymous):

6 divided 2/3 = N Find the value of N? My answer was 9, am I correct?

jhonyy9 (jhonyy9):

yes is right

OpenStudy (anonymous):

thank u

jhonyy9 (jhonyy9):

6/2/3 =18/2 =9